# Self-Normalized Resets for Plasticity in Continual Learning

### v2

@@ 1 Introduction
Plasticity Loss is an increasingly important phenomenon studied broadly under the rubric of continual learning (Dohare et al., 2024). This phenomenon refers to the empirical observation that as a network is continually trained on a sequence of changing tasks, its ability to adapt to a new task diminishes over time. While this is distinct from the problem of catastrophic forgetting (also studied under the rubric of continual learning (Goodfellow et al., 2013; Kirkpatrick et al., 2017)), it is of significant practical importance. In the context of pre-training language models, an approach that continually trains models with newly collected data is preferable to training from scratch (Ibrahim et al., 2024; Wu et al., 2024). On the other hand, the plasticity loss phenomenon demonstrates that such an approach will likely lead to models that are increasingly unable to adapt to new data. Similarly, in the context of reinforcement learning using algorithms like TD, where the learning tasks are inherently non-stationary, the plasticity loss phenomenon results in actor or critic networks that are increasingly unable to adapt to new data (Lyle et al., 2022). Figure 1 illustrates plasticity loss in the ‘Permuted MNIST’ problem introduced by Goodfellow et al. (2013).
@@
One formal definition of plasticity measures the ability of a network initialized at a specific set of parameters to fit a random target function using some pre-specified optimization procedure. In this sense, random parameter initializations (eg. Lyle et al. (2024)) are known to enjoy high plasticity. This has motivated two related classes of algorithms that attempt to mitigate plasticity loss. The first explicitly ‘resets’ neurons that are deemed to have low ‘utility’ (Dohare et al., 2023; Sokar et al., 2023). A reset re-initializes the neurons input weights and bias according to some suitable random initialization rule, and sets the output weights to zero; algorithms vary in how the utility of a neuron is defined and estimated from online data. A second class of algorithms perform this reset procedure implicitly via regularization (Ash & Adams, 2020; Kumar et al., 2023b). These latter algorithms differ in their choice of what to regularize towards, with choices including the original network initialization; a new randomly drawn initialization; or even zero. The aforementioned approaches to mitigating plasticity loss attempt to adjust the training process; other research has studied the role of architectural and optimizer hyperparameter choices. Across all of the approaches to mitigating plasticity loss described above, no single approach is yet to emerge as both robust to hyperparameter choices, and simultaneously performant across benchmark problems.
@@
Given some point process consider the task of distinguishing between the hypotheses that this point process has a positive rate (the null hypothesis), or a rate that is identically zero with a penalty for late rejection or acceptance. An optimal test here takes the following simple form: we reject the null hypothesis as soon as the time elapsed without an event exceeds some percentile of the inter-arrival time under the null hypothesis and otherwise accept immediately upon an event. Viewing the firing of a neuron as such a point process, we propose to reset a neuron based on a rejection of the hypothesis that the the neuron is firing at a positive rate. We use the histogram of past inter-firing times as a proxy of the inter-arrival time distribution under the null hypothesis. This exceedingly simple algorithm is specified by a single hyperparameter: the rejection percentile threshold. We refer to this procedure as self-normalized resets (SNR) and argue this is a promising approach to mitigating plasticity loss:
@@
    We demonstrate superior performance on four benchmark problems classes studied in (Dohare et al., 2023; Kumar et al., 2023b). Interestingly, there is no single closest competitor to SNR across these problems. Many competing approaches also show significant sensitivity to the choice of hyperparameters; SNR does not. We introduce a new problem to elucidate similar plasticity loss phenomena in the context of language models, and show similar relative merits for SNR.
@@
    We conduct a theoretical investigation of the optimization landscape for the problem of learning a single ReLU. We show that while (an idealized version of) SNR learns the target ReLU, regularization based approaches can fail to learn in this simple setting.
@@ 1.1 Related Literature
The phenomenon of plasticity loss was discovered in the context of transfer learning (Ash & Adams, 2020; Zilly et al., 2021; Achille et al., 2017). Achille et al. (2017) showed that pre-training a network on blurred CIFAR images reduces its ability to learn on the original images. In a similar vein, Ash & Adams (2020) showed that pre-training a network on 50% of a training set followed by training on the complete training set reduces accuracy relative to a network that forgoes the pre-training step. More recent literature has focused on problems that induce plasticity loss while training on a sequence of hundreds of changing tasks, such as Permuted MNIST and Continual ImageNet in Dohare et al. (2021), capturing the necessity to learn indefinitely.
@@
Regularization Methods. L2 regularization has been shown to reduce plasticity loss, but is insufficient in completely alleviating the phenomenon (Dohare et al., 2021; Lyle et al., 2023). While L2 regularization limits weight norm growth during continual learning, it can exacerbate weight rank collapse due to regularization towards the origin. One successful regularization technique is Shrink and Perturb (S&P) (Ash & Adams, 2020), which periodically scales the network’s weights by a shrinkage factor $p$ followed by adding random noise to each weight with scale $\sigma$. Another approach is to perform L2 regularization towards the initial weights referred to as L2 Init (Kumar et al., 2023b). These methods can be viewed as variants of L2 regularization that regularize towards a random initialization and the original initialization, respectively. These methods limit the growth of weight norms while maintaining weight rank and neuron activity by regularizing towards a high-plasticity parameterization.
@@ 2 Algorithm
One hypothesis that seeks to explain plasticity loss is that the network weights obtained from minimizing loss over some task yield poor initializations for a subsequent task, leading to the inactive neurons we observe in the above experiment. On the other hand random weight initializations are known to work well (Glorot & Bengio, 2010), suggesting a natural class of heuristics: re-initialize inactive neurons. Of course, the crux of any such algorithm is determining whether a neuron is inactive in the first place, and doing so as quickly as possible.
To motivate our algorithm, SNR, consider applying the network $f(\cdot;\theta_{t}^{*})$ to a hypothetical sequence of training examples drawn i.i.d. from $\mu_{t}$ indexed by $s$. Let $Z^{\mu_{t}}_{s,i}$ indicate the sequence of activations of neuron $i$, and let $A^{\mu_{t}}_{i}$ be a random variable distributed as the random time between any two consecutive activations over this hypothetical sequence of examples. Now turning to the actual sequence of training examples, let $a^{t}_{i}$ count the time since the last firing of neuron $i$ prior to time $t$. Our (idealized) proposal is then exceedingly simple: reset neuron $i$ at time $t$ iff $\mathbb{P}(A_{i}^{\mu_{t}}\geq a^{t}_{i})\leq\eta$ for some suitably small threshold $\eta$. We dub this algorithm Self-Normalized Resets and present it as Algorithm 1. The algorithm requires a single hyper parameter, $\eta$. Of course in practice, the distribution of $A_{i}^{\mu_{t}}$ is unknown to us, and so an implementable version of Algorithm 1 simply approximates this distribution with the histogram of inter-firing times of neuron $i$ prior to time $t$ [^3] As opposed to tracking the histogram itself, we simply track the mean inter-firing time, and assume $A_{i}^{\mu_{t}}$ is geometrically distributed with that mean. This requires tracking just one parameter per neuron. Our mean estimate is itself computed over a fixed length trailing window motivated by the change-point detection approach of Besbes & Zeevi (2011); the length of this window is a hyper-parameter.

### v3

@@ 1 Introduction
Plasticity Loss is an increasingly important phenomenon studied broadly under the rubric of continual learning (Dohare et al., 2024). This phenomenon refers to the empirical observation that as a neural network is continually trained on a sequence of changing tasks, its ability to adapt to a new task diminishes over time. While this is distinct from the problem of catastrophic forgetting (also studied under the rubric of continual learning (Goodfellow et al., 2013; Kirkpatrick et al., 2017)), it is of significant practical importance. In the context of pre-training language models, an approach that continually trains models with newly collected data is preferable to training from scratch (Ibrahim et al., 2024; Wu et al., 2024). On the other hand, the plasticity loss phenomenon demonstrates that such an approach will likely lead to models that are increasingly unable to adapt to new data. Similarly, in the context of reinforcement learning using algorithms like TD, where the learning tasks are inherently non-stationary, the plasticity loss phenomenon results in actor or critic networks that are increasingly unable to adapt to new data (Lyle et al., 2022). Figure 1 illustrates plasticity loss in the ‘Permuted MNIST’ problem introduced by Goodfellow et al. (2013).
@@
One formal definition of plasticity measures the ability of a network initialized at a specific set of parameters to fit a random target function using some pre-specified optimization procedure. In this sense, random parameter initializations (eg. Lyle et al. (2024)) are known to enjoy high plasticity. This has motivated two related classes of algorithms that attempt to mitigate plasticity loss. The first explicitly ‘resets’ neuron’s that are deemed to have low ‘utility’ (Dohare et al., 2023; Sokar et al., 2023). A reset re-initializes the neurons input weights and bias according to some suitable random initialization rule, and sets the output weights to zero; algorithms vary in how the utility of a neuron is defined and estimated from online data. A second class of algorithms perform this reset procedure implicitly via regularization (Ash and Adams, 2020; Kumar et al., 2023b). These latter algorithms differ in their choice of what to regularize towards, with choices including the original network initialization; a new randomly drawn initialization; or even zero. The aforementioned approaches to mitigating plasticity loss attempt to adjust the training process; other research has studied the role of architectural and optimizer hyperparameter choices. Across all of the approaches to mitigating plasticity loss described above, no single approach is yet to emerge as both robust to hyperparameter choices, and simultaneously performant across benchmark problems.
Several correlates of plasticity loss have been identified such as neuron inactivity, feature or weight–rank collapse, increasing weight norms, and loss of curvature in the loss surface (Dohare et al., 2021; Lyle et al., 2023; Sokar et al., 2023; Lewandowski et al., 2023; Kumar et al., 2020). Despite the extensive documentation of the phenomenon, the literature remains almost entirely empirical, offering little theoretical insight into why plasticity fades or how best to restore it. To address this gap, we analyze the simplest non-trivial continual learning model: a single ReLU-activated neuron trained by gradient descent on adversarially selected targets. We show that gradient descent, even with $L_{2}$ regularization, incurs non-vanishing average regret, whereas gradient descent augmented with an idealized reset-oracle achieves $\mathcal{O}(d^{2}\ln{T}/T)$ average regret. Although the problem of learning a single ReLU-activated neuron has been studied extensively, no prior work considers the case where targets are chosen adversarially with knowledge of the initialization. Our paper is the first to analyze this problem under this adversarial regime.
Having thus characterized plasticity loss for the fundamental primitive of a single ReLU-activated neuron, we turn to the practical question of how to detect when a neuron has effectively “died” as training data arrives sequentially. To this end, we cast the problem of detecting a dead neuron as an optimal hypothesis‐test problem. Specifically, when minimizing type-1 and type-2 error and a delay penalty of detecting a dead neuron, we show that the optimal reset criterion is to define a reset-threshold that is proportional to the neuron’s nominal firing rate $p$. This optimal solution is precisely our self-normalized resets (SNR) algorithm. In contrast to SNR, existing reset-criteria define a single reset-threshold or reset-frequency for a neural network. We make evident the efficacy of SNR in relation to such reset-criteria by showing that for the problem of detecting neuron death of two neurons with nominal firing rates $p_{1}<p_{2}$, the error rate under an (optimal) fixed threshold grows arbitrarily larger than the error rate under SNR as (a) the penalty for delays shrinks to zero or (b) the gap between the nominal firing rates $p_{1}$ and $p_{2}$ widens. Concretely, we introduce the SNR algorithm and distill the key intuition behind it as follows:
@@
Given some point process consider the task of distinguishing between the hypotheses that this point process has a positive rate (the null hypothesis), or a rate that is identically zero with a penalty for late rejection or acceptance. An optimal test here takes the following simple form: we reject the null hypothesis as soon as the time elapsed without an event exceeds some percentile of the inter-arrival time under the null hypothesis and otherwise accept immediately upon an event. Viewing the firing of a neuron as such a point process, we propose to reset a neuron based on a rejection of the hypothesis that the the neuron is firing at a positive rate. We use the histogram of past inter-firing times as a proxy of the inter-arrival time distribution under the null hypothesis. This exceedingly simple algorithm is specified by a single hyperparameter: the rejection percentile threshold. We refer to this procedure as self-normalized resets (SNR) and argue this is a promising approach to mitigating plasticity loss.
1.1 Contributions
With the objective of studying the phenomenon of plasticity loss from a theoretical standpoint and thus motivating an efficient remedy, we make the following contributions.
@@ 1.1 Contributions
    The Self-Normalized Resets Algorithm. We introduce the Self-Normalized Resets (SNR) algorithm and argue its efficacy from both a theoretical and empirical standpoint.
@@
    Characterizing Plasticity Loss and the Efficacy of Neuron-Resets. We conduct a theoretical investigation of the optimization landscape for the problem of learning a single RELU-activated neuron with gradient descent. While existing literature has solely considered the case of non-adversarial target weights in which the target weights are first fixed and then the network weights are sampled, we introduce the setting of adversarial target weight selection in which the network weights are first sampled and then the target weights are selected adversarially with knowledge of the network weight initialization. Crucially, this setting captures the essence of continual learning, whereby consecutive learning tasks can be unrelated or even adversarially selected. Under this setting, we show in Theorem 3.2 that gradient descent, even with $L_{2}$ regularization, attains non-vanishing average regret. In addition, in Theorem 3.3 we show that gradient descent with access to an idealized reset-oracle attains $O(d^{2}\ln{T}/T)$ expected average regret. Our theoretical analysis characterizes plasticity loss, for the simple primitive of a single neuron, as being a consequence of network weights learned for one task yielding a poor local minimum for a subsequent task. Moreover, our analysis establishes neuron resets to be an effective remedy to plasticity loss while $L_{2}$-regularization is insufficient.
  3. 3.
    Deriving an Optimal Neuron-Reset Mechanism. To answer the question of what is an effective reset scheme, we model the problem of detecting neuron death as that of an optimal hypothesis test with the objective of minimizing type-1 and type-2 error and a delay penalty (scaled by $\lambda$). In Proposition 3.1 we show that for a neuron with a nominal firing rate $p>0$, the optimal reset threshold is the $1-\lambda(p-\lambda)^{-1}$ percentile of a Geometric(p)(p) distribution; this is precisely the SNR reset heuristic. We juxtapose the SNR reset criterion with existing neuron-reset schemes, which define a fixed reset-threshold or reset-frequency $r^{*}$, by considering the problem of detecting neuron death of two neurons with nominal firing rates $p_{1}<p_{2}$. Specifically, in Proposition 3.2 we show that the total error incurred by a fixed threshold $r^{*}$ to that under SNR scales like
    where $\alpha$ relates the delay penalty factors according to $\lambda_{1}=\alpha p_{1}$ and $\lambda_{2}=\alpha p_{2}$.
  4. 4.
    Empirically Validating SNR. We demonstrate superior performance on four benchmark problems studied in (Dohare et al., 2023; Kumar et al., 2023b) using MLP, CNN, and ViT architectures. Interestingly, there is no single closest competitor to SNR across these problems. Many competing approaches also show significant sensitivity to the choice of hyperparameters; SNR does not. Additionally we introduce a new problem, Permuted Shakespeare, to elucidate similar plasticity loss phenomena in the context of language models, and show similar relative merits for SNR.
@@ 1.2 Related Literature
The phenomenon of plasticity loss was discovered in the context of transfer learning (Ash and Adams, 2020; Zilly et al., 2021; Achille et al., 2017). Achille et al. (2017) showed that pre-training a network on blurred CIFAR images reduces its ability to learn on the original images. In a similar vein, Ash and Adams (2020) showed that pre-training a network on 50% of a training set followed by training on the complete training set reduces accuracy relative to a network that forgoes the pre-training step. More recent literature has focused on problems that induce plasticity loss while training on a sequence of hundreds of changing tasks, such as Permuted MNIST and Continual ImageNet in Dohare et al. (2021), capturing the necessity to learn indefinitely.
@@
Regularization Methods. L2 regularization has been shown to reduce plasticity loss, but is insufficient in completely alleviating the phenomenon (Dohare et al., 2021; Lyle et al., 2023). While L2 regularization limits weight norm growth during continual learning, it can exacerbate weight rank collapse due to regularization towards the origin. One successful regularization technique is Shrink and Perturb (S&P) (Ash and Adams, 2020), which periodically scales the network’s weights by a shrinkage factor $p$ followed by adding random noise to each weight with scale $\sigma$. Another approach is to perform L2 regularization towards the initial weights referred to as L2 Init (Kumar et al., 2023b). These methods can be viewed as variants of L2 regularization that regularize towards a random initialization and the original initialization, respectively. These methods limit the growth of weight norms while maintaining weight rank and neuron activity by regularizing towards a high-plasticity parameterization.
@@
Learning a single ReLU-Activated Neuron with Gradient Descent. Learning a single ReLU-activated neuron with gradient-based methods is a fundamental problem that has been extensively studied in the theory literature. Early works analyzed the bias-free case, for example, Yehudai and Ohad (2020) showed that with a well-spread input distribution (e.g. uniform over a sphere), gradient methods can recover the target weight vector with high probability. Vardi et al. (2021) extended this line of work to a neuron with bias, finding that the addition of a bias significantly complicates the optimization landscape. Nevertheless, under appropriate assumptions on the data distribution (such as isotropic or sufficiently full support in every direction) and on the initialization, they proved that gradient descent will still converge to the global minimum, with high probability at a linear rate. Beyond single neurons, more general models like max-affine regression (which includes learning a ReLU as a special case) have also been studied – Kim and Lee (2024) show that first-order methods can achieve linear convergence for such piecewise-linear regression problems, assuming i.i.d. sub-Gaussian inputs and a suitable initialization. Importantly, all these results assume the target neuron’s parameters are fixed a priori (non-adversarial) and the data are drawn from a benign distribution. To our knowledge, no prior work addresses the setting of adversarially selected target weights – this paper is the first to analyze gradient descent for learning a ReLU neuron in an adversarial target regime, where the ground-truth weight vector may be chosen by an adversary (with knowledge of the initialization).
1.3 Roadmap of the Paper
We outline the roadmap for the rest of this paper.
In Section 2 we formally define plasticity loss as an empirical phenomenon that corresponds to an increasing average expected regret over a stream of training data. In addition, we formally define the Self-Normalized Resets algorithms and present a simple benchmark problem, Permuted MNIST, that illustrates the phenomenon of plasticity loss and the efficacy of SNR.
In Section 3 we present our theoretical analysis. In Section 3.1.1 we formally introduce the problem of learning a single ReLU-activated neuron. In Section 3.1.3 we formally define the setting of adversarial target weights and present Theorem 3.2. In Section 3.1.4 we introduce the idealized reset oracle, define the gradient descent algorithm augmented with such a reset oracle, and present Theorem 3.3. In Section 3.2 we formally introduce the hypothesis test problem, that captures the problem of neuron-death detection, and we present Proposition 3.1 and Proposition 3.2.
In Section 4 we present our empirical results demonstrating the efficacy of SNR across a series of continual learning benchmarks and architectures. In addition, we discuss pertinent phenomena that we observe in our experiments.
Finally, in Section 5 we conclude with a Discussion and Limitations section, summarizing the our key contributions and noting limitations.
@@ 2 Algorithm
One hypothesis that seeks to explain plasticity loss is that the network weights obtained from minimizing loss over some task yield poor initializations for a subsequent task, leading to the inactive neurons we observe in the above experiment. On the other hand random weight initializations are known to work well (Glorot and Bengio, 2010), suggesting a natural class of heuristics: re-initialize inactive neurons. Of course, the crux of any such algorithm is determining whether a neuron is inactive in the first place, and doing so as quickly as possible.
To motivate our algorithm, SNR, consider applying the network $f(\cdot;\theta_{t}^{*})$ to a hypothetical sequence of training examples drawn i.i.d. from $\mu_{t}$ indexed by $s$. Let $Z^{\mu_{t}}_{s,i}$ indicate the sequence of activations of neuron $i$, and let $A^{\mu_{t}}_{i}$ be a random variable distributed as the random time between any two consecutive activations over this hypothetical sequence of examples. Now turning to the actual sequence of training examples, let $a^{t}_{i}$ count the time since the last firing of neuron $i$ prior to time $t$. Our (idealized) proposal is then exceedingly simple: reset neuron $i$ at time $t$ iff $\mathbb{P}(A_{i}^{\mu_{t}}\geq a^{t}_{i})\leq\eta$ for some suitably small threshold $\eta$. We dub this algorithm Self-Normalized Resets and present it as Algorithm 1. The algorithm requires a single hyper parameter, $\eta$. Of course in practice, the distribution of $A_{i}^{\mu_{t}}$ is unknown to us, and so an implementable version of Algorithm 1 simply approximates this distribution with the histogram of inter-firing times of neuron $i$ prior to time tt. [^3] As opposed to tracking the histogram itself, we simply track the mean inter-firing time, and assume $A_{i}^{\mu_{t}}$ is geometrically distributed with that mean. This requires tracking just one parameter per neuron. Our mean estimate is itself computed over a fixed length trailing window motivated by the change-point detection approach of Besbes and Zeevi (2011); the length of this window is a hyper-parameter.
@@
Algorithm 1 SNR: Self-Normalized Resets
3 Theoretical Analysis
In this section we provide a two-fold theoretical analysis. We begin in Section 3.1 by analyzing the problem of learning a single ReLU-activated neuron with gradient descent under an adversarial selection of the target weights. This is the simplest possible (component of a) neural network and it captures the essence of continual learning, whereby the adversarial selection of the target weights models task transitions. We begin by showing that gradient descent with L2 regularization attains $\Omega(1)$ expected average regret. This analysis characterizes plasticity loss, at the neuron-level, as being a consequence of network weights learned for previous tasks forming poor initializations for subsequent tasks. Secondly, we provide a positive result demonstrating that gradient descent with access to a reset oracle attains $O(d^{2}\ln{T}/T)$ expected average regret, demonstrating that neuron resets are an effective mechanism for identifying poor initializations and thus remedying plasticity loss.
Having characterized plasticity loss and shown that resets are an effective remedy in Section 3.1, in Section 3.2 we answer the question of what makes an effective reset oracle in practice when data is arriving sequentially. To this end, we frame the problem of detecting an inactive neuron as that of an optimal hypothesis test. We show that an optimal reset scheme must have a reset threshold that is proportional to the neuron’s firing rate; this is precisely the SNR mechanism. In contrast, existing competitor reset schemes define a fixed reset threshold or frequency. Our experiments in Section 4 show that such reset schemes attain performance inferior to that of SNR. To shed intuition on this observation, we show that under this optimal hypothesis test model, one can incur arbitrarily large error relative to SNR if attempting to detect two dead neurons with distinct firing rates with a single threshold.
3.1 Characterizing Plasticity Loss in a Single ReLU and the Promise of Resets
In this section, we provide a theoretical analysis of learning a single ReLU-activated neuron with gradient descent and with access to a reset oracle, where the target neuron’s weights are chosen adversarially after the network’s weight initialization is fixed. This adversarial selection serves as an abstraction of task transitions in a continual learning setting, allowing for scenarios where successive tasks may be unrelated—or even deliberately chosen to be adversarial. In contrast, prior work has solely considered settings where the target neuron’s weights are fixed before the network’s initialization is sampled (Vardi et al., 2021; Kim and Lee, 2024; Yehudai and Ohad, 2020). While such literature demonstrate the effectiveness of first-order methods for learning a single neuron and the benefits of modern weight initializations, they do not capture the challenges of task transitions, where the initialization for a new task may be influenced by prior learning.
3.1.1 Preliminaries
To study the learning dynamics of a single ReLU-activated neuron, we consider the expected squared error:
$L(w)=\mathbb{E}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}],$ (1)
where $\sigma(z)=\max(0,z)$ is the ReLU activation function, $v$ is the target neuron’s weight vector, and the expectation is taken over $x\in\mathbb{R}^{d+1}$, sampled from $\mathcal{N}(0,I_{d})\times\{1\}$, where $I_{d}$ is the $d$-dimensional identity matrix. The gradient of the loss is then given by
$\nabla L(w)=\mathbb{E}[2(\sigma(w^{\top}x)-\sigma(v^{\top}x))x\mathbbm{1}_{\{w^{\top}x\geq 0\}}].$
where we define the subgradient of $\sigma$ as $\sigma^{\prime}(z)=\mathbbm{1}_{\{z\geq 0\}}$, which corresponds to a particular choice of subgradient at $z=0$. Next, we introduce the $L_{2}$-regularized loss function $\hat{L}(w)$, given by
$\hat{L}(w)=\mathbb{E}\bigl[(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}\bigr]+\frac{\lambda}{2}\|w\|^{2},$ (2)
where $\lambda\geq 0$ is the regularization parameter. Setting $\lambda=0$ recovers the original loss function $L(w)$. The additional $L_{2}$ regularization term contributes an additive term $\lambda w$ to the gradient, yielding
$\nabla\hat{L}(w)=\nabla L(w)+\lambda w=\mathbb{E}\bigl[2(\sigma(w^{\top}x)-\sigma(v^{\top}x))x\mathbbm{1}_{\{w^{\top}x\geq 0\}}\bigr]+\lambda w.$
For an iterate $w_{t}$, step size $\alpha$, and a general loss function $\mathcal{L}(w)$, which may represent either the standard loss $L(w)$ or the $L_{2}$-regularized loss $\hat{L}(w)$, the gradient descent update is defined as
$w_{t+1}=w_{t}-\alpha\nabla\mathcal{L}(w_{t}).$
Finally, we define the average regret in learning the target neuron’s weights as
$R_{T}=\frac{1}{T}\sum_{t=0}^{T-1}\left(L(w_{t})-L(w^{*})\right)=\frac{1}{T}\sum_{t=0}^{T-1}L(w_{t}),$ (3)
where the second equality follows from the fact that the optimal weight vector for (1) is $w^{*}=v$, and thus
$L(w^{*})=L(v)=0.$
We also define the expected average regret, $\mathbb{E}[R_{T}]$, which accounts for randomness in the learning process. Specifically, the expectation is taken over any stochasticity introduced by the learning algorithm, including the initialization of $w_{0}$ and any reinitializations of $w_{t}$ due to the reset oracle. Thus, the goal of learning a ReLU-activated neuron is to minimize (expected) average regret.
3.1.2 Non-Adversarial Target-Weights
Prior work on gradient‐descent dynamics has almost exclusively considered the regime in which the target weights $v$ are fixed before the network weights $w_{0}$ are drawn from a prescribed initialization distribution Vardi et al. (2021); Yehudai and Ohad (2020); Kim and Lee (2024). For clarity, Algorithm 2 formalizes gradient descent under this non-adversarial setup.
Target Selection : Target weights $v$ are fixed
Weight Initialization : Initialize $w_{0}$ from some distribution
for $t\leftarrow 0,1,2,\dots$ do
   Compute gradient $\nabla\mathcal{L}(w_{t})$
   Update weights: $w_{t+1}\leftarrow w_{t}-\alpha\,\nabla\mathcal{L}(w_{t})$